OpenStudy (anonymous):

The gas tank of a car is 36% empty. What is the fraction of the tank is full? A) 36/100 B) 16/25 C) 12/25 D) 9/25

OpenStudy (anonymous):

100-36= 64 now 64/100= 32/50= 16/25 so B

OpenStudy (anonymous):

Oh okay. Thank you!

OpenStudy (anonymous):

np!

OpenStudy (anonymous):

all you had o do was the steps that I said then your good to go!